OpenStudy (anonymous):

Solve the system of equations using the elimination method. Enter the answer as an ordered pair. 3y = x - 1 x - 2y = 2

OpenStudy (anonymous):

-x+3y =-1 (Rearrange the first equation) x - 2y = 2 Then add the two equations together and you get: y=1 Plug y=1 back into one of the equation and solve for x. x-2y=2 x=2y+2 x=2*1+2 x=2+2 x=4 (4,1) is the answer

OpenStudy (anonymous):

@stargirlzilin Thankyou